pearis:

PleaseHelp!! A Ferris wheel car moves from point C to point D on the circle shown below: Circle A is shown with a central angle marked 37 degrees and the diameter marked 35 feet. What is the arc length the car traveled, to the nearest hundredth? (5 points)

pearis:

so do I solve for that equation? cause here are the answer choices 8.30 feet 11.30 feet 12.41 feet 13.15 feet

pearis:

hold on let me solve it

justus:

Do you still need help?

AngeI:

Hello welcome to QuestionCove, may you pleas attach the circle that is supposed to be used?

jhonyy9:

A Ferris wheel car moves from point C to point D on the circle shown below: Circle A is shown with a central angle marked 37 degrees and the diameter marked 35 feet. What is the arc length the car traveled, to the nearest hundredth?  so something in this way ? than any idea now ? @XioGonz

XioGonz:

Okay Formula: s = (x/360)*pi*d s = arc length = unknown x = central angle = 37 degrees d = diameter So= 35 ft s = (x/360)*pi*d Plug in the given numbers s = (37/360)*pi*35 s = 11.30 SO your answer is Choice B) 11.30 feet